St. Blues Technologies' expected (next year) EBIT is $292.00, its tax rate is 40%, depreciation is...

  1. St. Blues Technologies' expected (next year) EBIT is $292.00, its tax rate is 40%, depreciation is $18.00, planned capital expenditures are $80.00, and planned INCREASES in net working capital is $24.00.

What is the free cash flow to the firm (FCFF)?

$

The firm's interest expense is $24.00. Assume the tax rate is 40% and the net debt of the firm DECREASES by $5.00.

What is the free cash flow to equity (FCFE)?

$

What is the market value of equity if the FCFE is projected to grow at 4% indefinitely and the cost of equity is 11%?  (Round this answer to 2 decimal places.)

$
Answer #1

Given for St. Blues Technologies

EBIT = $292

Tax rate = 40%

Depreciation = $18

Planned capex = $80

&  planned INCREASES in net working capital = $24.00.

1). Free cash flow to firm FCFF = EBIT*(1-T) + Depreciation - Planned Capex - Change in Working capital

FCFF = 292*0.6 + 18 - 80 - 24 = $89.20

2). Interest expenses = $24

Decrease in net borrowing = $5.00

Free cash flow to equity(FCFE) = FCFF – Int(1 – Tax rate) + increase Net borrowing.

FCFE = 89.20 - 24*0.6 - 5 = $69.80

3). FCFE is projected to grow at 4% indefinitely.

Cost of equity Ke = 11%

Market value of equity can be calculated by summing up present value of all the future FCFE at a discount rate of Ke.

Using perpetuity formula

Market value of equity = FCFF1/(Ke-g) = 69.80/(0.11-0.04) = $997.14